# leftmargin=12.5pt

### v1

@@ I Introduction
We know that the non-relativistic limit of the relativistic physics plays an important role in modern physics. These non-relativistic theories have found applications in diverse fields [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. The non-relativistic limit can be obtained by group contraction method from Poincare group which describes the relativistic sector. One of the well studied examples of non-relativistic limit is the Galilean limit where one considers the speed of light to be considerably greater than the other speeds. The study of Galilean relativistic theories was initiated by Le Bellac and Levy Leblond [11] back in 1970’s. Afterwards various aspects of Galilean relativistic physics have been studied [12, 13, 14, 15, 16].
@@
Here in this paper we are mostly interested in the formulation aspect of the Carroll relativistic field theory. Almost all the work in this field follow the formalism put forward by Levy-Leblond [17] which is based on the argument of the principle of causality. We, on the other hand are more interested by the kinematic approach proposed by N. D. Sengupta [18]. He gave a new set of transformations from which we are able to derive known Carroll relativistic results. However Sengupta’s work [18] received little recognition  [^2] Very recently Levy-Leblond quite decently mentioned about Sengupta’s work in a review [30]. Our main motivation here is to build up a systematic action formalism for a Carroll relativistic field theory following the similar track that we have used for the Galilean relativistic field theories [14, 15, 16]. One cannot get the desired results by starting with usual Lorentz transformations simply because they are not valid. Here Sengupta’s formulation plays a crucial role. We show in our paper in a very straight forward way how to build up a consistent Carroll relativistic action formulation by using Sengupta’s formulation [18].
@@ VI Transformations of fields under boost
It is interesting to note that the scaling relations for the fields may be derived directly using boost transformations. The appropriate boost transformations are reproduced for the scaling relations given in table 2. This provides an additional justification for taking those relations. Now to proceed further one has to compare the module of the electric field with that of the magnetic field. So for the electric limit, where electric field is dominating than the magnetic field, we get

### v2

# Sengupta Transformations and Carrollian Relativistic Theory

@@ I Introduction
We know that the non-relativistic limit of the relativistic physics plays an important role in modern physics. These non-relativistic theories have found applications in diverse fields [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. The non-relativistic limit can be obtained by group contraction method from Poincare group which describes the relativistic sector. One of the well studied examples of non-relativistic limit is the Galilean limit where one considers the speed of light to be considerably greater than the other speeds. The study of Galilean relativistic theories was initiated by Le Bellac and Levy Leblond [11] back in 1970’s. Afterwards various aspects of Galilean relativistic physics have been studied [12, 13, 14, 15, 16]. A very good discussion on the physical implications of Galilean electrodynamics can be found in the review article [17]. [^2] Galilean aspect of Coulomb and Lorentz gauge has been discussed in [40] and the usage of Galilean equations in the context of momentum transfer is shown in [41].
@@
Here in this paper we are mostly interested in the formulation aspect of the Carroll relativistic field theory. Almost all the work in this field follow the formalism put forward by Levy-Leblond [18] which is based on the argument of the principle of causality. For example a Hamiltonian formulation of Carrollian relativistic theory is provided in [31]. We, on the other hand are more interested by the kinematic approach proposed by N. D. Sengupta [19]. He gave a new set of transformations from which we are able to derive known Carroll relativistic results. However Sengupta’s work [19] received little recognition  [^3] Very recently Levy-Leblond quite decently mentioned about Sengupta’s work in a review [32]. Our main motivation here is to build up a systematic action formalism for a Carroll relativistic field theory following the similar track that we have used for the Galilean relativistic field theories [14, 15, 16] using Sengupta’s transformations. One cannot get the desired results by starting with usual Lorentz transformations simply because they are not valid. Here Sengupta’s formulation plays a crucial role. We show in our paper in a very straight forward way how to build up a consistent Carroll relativistic action formulation by using Sengupta’s formulation [19].
@@ VI Transformations of fields under boost
It is interesting to note that the scaling relations for the fields may be derived directly using boost transformations. For the Galilean electrodynamics a similar exercise has been done in [39]. The appropriate boost transformations are reproduced for the scaling relations given in table 2. This provides an additional justification for taking those relations. Now to proceed further one has to compare the module of the electric field with that of the magnetic field. So for the electric limit, where electric field is dominating than the magnetic field, we get
@@ IX Conclusions
Carrollian electrodynamics has been discussed earlier as well. However our approach is different from these attempts not only from the construction point of view, but also in terms of the robustness of the physical arguments. In [13], the same has been constructed through a geometrical point of view. However it did not introduce the action in the Carrollian limit. [36] also briefly mentioned about the same in the light of flat space holography, however without any systematic analysis. Their discussion does not provide the physical construction of the underlying scaling nature of the potential vector components and therefore carries a possible ambiguity in the analysis. Although [37] captures the action formulation of the Carroll electrodynamics, but their approach to construct the same is different. We have started from Sengupta transformation to motivate the scaling behaviour of the Carroll vectors while [37] considered the Lorentz transformations as the guiding one. They have expanded the Lorentz transformation in powers of $c$ to obtain the required scaling laws. A similar expansion method on the potential vector has been adopted in [38] to derive the Carroll limit of Yang-Mills theory. In contrary the Sengupta transformations played the pivotal role in present analysis and it helped to obtain a systematic construction staring from the action formulation of Carrollian electrodynamics. In this sense it provides a much robust and physically motivated formulation. Galilean transformations are non-relativistic limit of Lorentz transformations. Analogously we derived here the Carroll transformations from the Sengupta ones.